The distributions with large n don't need to be normal, but these are.

In this problem you do not know the population standard deviation.

The mean for the first set ¯xx¯ 1 = 12.346 with a standard deviation of s1 = 2.215 There sample size was 14.

The mean for the first set ¯xx¯  2  = 12.181 with a standard deviation of s2 = 1.888 There sample size was 21.

Find the degrees of freedom.   (round to 2 places) The formula is the messy one on the MOM chapter 10 instruction page.

Use a two tail test. Find the left tail critical value, t-star for alpha = 0.01. Use the truncated version of the degrees of freedom.   (Remember the right tail is just the positive version. Use the inverse of t on excel. You will get a negative number.)

Find the test statistic t=   Round to 4 places.
Answer #1

Answer:

Given Data

\bar{x}_{1} = 12.346

s1 = 2.215

11 = 14

\bar{x}_{2}. = 12.181

s2 = 1.888

2 =21.

We assume that unequal variances.

1) The degrees of freedom

  df=\frac{\left [ \frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}} \right ]^{2}}{\frac{\left ( \frac{s_{1}^{2} }{n_{1}}\right )^{2}}{n_{1}-1}+\frac{\left ( \frac{s_{2}^{2} }{n_{2}}\right )^{2}}{n_{2}-1}}

=\frac{\frac{(2.25)^{2}}{14}+\frac{(1.888)^{2}}{21}}{\frac{\left ( \frac{5.0625}{14} \right )^{2}}{14-1}+\frac{\left ( \frac{(3.5645)}{21} \right )^{2}}{21-1}}

= \frac{0.36161+0.1697}{0.0101+0.00144}

= \frac{0.53131}{0.01154}

= 46.0407279

  df\cong 46

2)  \alpha = 0.01

It is two tailed test .

\alpha /2= 0.005

t_{\alpha/2,df}=t_{0.005,46}

= \pm 2.9487

The Right critical value = + 2.9487

3) The test Statistic:

t=\frac{\bar{X}_{1}-\bar{X_{2}}}{\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}}

=\frac{12.346-12.181}{\sqrt{\frac{4.906225}{14}+\frac{3.564544}{21}}}

  =\frac{0.165}{\sqrt{0.35044+0.16974}}

=\frac{0.165}{0.72124}

= 0.22877

****Please like it....

It is important to me....

Thank you for supporting me...